OpenStudy (anonymous):

3. There are two fruit trees located at (3,0) and (–3, 0) in the backyard plan. Maurice wants to use these two fruit trees as the focal points for an elliptical flowerbed. Johanna wants to use these two fruit trees as the focal points for some hyperbolic flowerbeds. Create the location of two vertices on the y-axis. Show your work creating the equations for both the horizontal ellipse and the horizontal hyperbola. Include the graph of both equations and the focal points on the same coordinate plane.

OpenStudy (anonymous):

This is my last question on my last assignment and I have no clue on how to answer it

OpenStudy (anonymous):

In which grade u are ?

OpenStudy (anonymous):

Finishing 10th

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

 would this work as the ellipse part of the question?

OpenStudy (zale101):

i think what u did is correct

OpenStudy (anonymous):

confusing question lol

OpenStudy (zale101):

they want to draw an equation for both ellipse and hyperbola, right?

OpenStudy (anonymous):

yea, i only did ellipse so far though

OpenStudy (anonymous):

i don't think he's there

OpenStudy (zale101):

the center is 0,0 horizontal ellipse equation is \[\frac{ (x-h)^2 }{ a^2 } +\frac{ (y-k)^2 }{ b^2 }=1\]